# Penalty decomposition derivative free method for the minimization of partially separable functions over a convex feasible set

### v1

@@ Abstract
In this paper, we consider the problem of minimizing a smooth function, given as finite sum of black-box functions, over a convex set. In order to advantageously exploit the structure of the problem, for instance when the terms of the objective functions are partially separable, noisy, costly or with first-order information partially accessible, we propose a framework where the penalty decomposition approach is combined with a derivative-free line search-based method. Under standard assumptions, we state theoretical results showing that the proposed algorithm is well-defined and globally convergent to stationary points. The results of preliminary numerical experiments, performed on test problems with number of variables up to thousands, show the validity of the proposed method compared with a standard derivative-free line search algorithm. Moreover, it is shown that the method is easily parallelizable and hence capable of taking advantage of parallelization of computation, when possible.
@@
Penalty decomposition, Finite-sum, Coordinate partially separable, Derivative-free line search
@@ 1 Introduction
$\begin{array}[]{l}\min f(x)=\displaystyle{\sum_{j=1}^{m}}f_{j}(x)\\ \text{s.t. }x\in X\end{array}$ (1)
where $X\subset\mathbb{R}^{n}$ is a closed, convex set and $f_{j}:\mathbb{R}^{n}\to\mathbb{R}$, for $j=1,\ldots,m$ are smooth functions for which, however, we do not have access to first-order information. We also assume that $f$ is bounded below by some value $f^{*}$.
The literature on derivative-free methods is wide and increasing. For insights on general derivative-free approaches we refer the reader to the books [5, 2], the survey [15] and to some seminal papers both for smooth optimization (e.g., [7, 23, 16, 22]) and nonsmooth optimization (e.g., [1, 8, 17]).
@@
Briefly, the penalty decomposition approach consists in the introduction of an auxiliary vector of variables - copies in principle of the original ones - and the addition of a penalty term to the objective function, penalizing the distance between each pair of corresponding variables. The introduction of the auxiliary variables allows the decomposition of the problem into smaller independent subproblems; this may be advantageously exploited to take into account the peculiarities of problem (1) in the cases (i)-(iv) discussed above.
@@
In this work we define a general algorithm model where the penalty decomposition strategy is coupled with a derivative-free line search-based method for the management of subproblems. We state theoretical results - not immediately descending from known results in the literature - for proving global convergence properties of the proposed algorithm and we present its characterization in the specific case of partially separable functions. Finally, we report the results of preliminary computational experiments.
@@ 2 A Penalty decomposition approach
A sequential penalty method [3, Ch. 4.2] can then be employed to address problem (3) and thus (2), which iteratively solves (inexactly) subproblems of the form
@@
[…]
for a sequence of values $\{\tau_{k}\}$ such that $\tau_{k}\to\infty$. The subproblem is considered solved if a solution $(x^{k},y^{k},z^{k})$ is found such that
@@
for some positive tolerance $\xi_{k}>0$ and constants $c_{y},c_{z}>0$. The resulting algorithm is formally described by Algorithm 1. Note that, formally, the solution $(x^{k-1},y^{k-1},z^{k-1})$ is not used during iteration $k$; however, it is well-known that, with sequential penalty type methods, using the solution of the previous subproblem as a starting point for the next one is often good for efficiency and stability reasons [3, Ch. 4.2]. We now turn to the convergence analysis of the sequential penalty framework.
1 Input: $x^{0},y^{0},z^{0}\in\mathbb{R}^{n},$$\{\xi_{k}\}\subseteq\mathbb{R}_{+},\{\tau_{k}\}\subseteq\mathbb{R}_{+}$, $c_{y},c_{z}>0$
@@ 2.1 Convergence analysis
We start the analysis stating the following reasonable assumption (formally given for the case of $f$ defined as the sum of $m$ terms).
@@
=$\mathcal{L}=\{(x,y,z)\in X\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mid P_{\tau}(x,y,z)\leq P_{\tau}(\bar{x},\bar{y},\bar{z})\}$
@@
We now make a second, reasonable assumption on the algorithm.
Assumption 2.
For all $k$, $(x^{k},y^{k},z^{k})$ satisfies
$P_{\tau_{k}}(x^{k},y^{k},z^{k})\leq P_{\tau_{0}}(x^{0},y^{0},z^{0}).$ (8)
Remark 1.
Note that, for any $\tau>0$ and any feasible $(x,y,z)$, we have
$P_{\tau}(x,y,z)=f(x).$
Therefore, if we start with a solution $(x^{0},y^{0},z^{0})$ such that $x^{0}\in X$ and $y^{0}=z^{0}=x^{0}$, condition (8) can certainly be satisfied if at iteration $k$ we generate $(x^{k},y^{k},z^{k})$ by means of a monotone descent method starting at
@@
Let $\{(x^{k},y^{k},z^{k})\}$ be a sequence satisfying Assumption 2 . Then $\{(x^{k},y^{k},z^{k})\}$ is a bounded sequence.
@@
is compact. Assumption 2 implies that
@@
The result in Proposition 2.4 cannot be directly inferred, to the best of our knowledge, by known results in the literature. Indeed, most known results concerning penalty and augmented Lagrangian methods rely on different stationarity conditions for subproblems with lower-level constraints to be handled explicitly out of the penalty function (see, e.g., [4]). On the other hand, the case of projection-based stationarity has been previously considered in [10, 9], but a bond between the sequences $\{\tau_{k}\}$ and $\{\xi_{k}\}$ is required ($\tau_{k}\xi_{k}\to 0$). For this particular setup we thus provide a novel, improved result of convergence that does not need this particular interaction between the two sequences of parameters.
@@ 3 A derivative-free alternate minimization algorithm for penalty subproblems
    fixed $y$ and $z$, we have a problem in $x$ whose globally optimal solution can be determined, being it convex quadratic with a separable objective function subject to convex constraints; the solution is in fact available in closed form in the particular case of bound constraints.
@@ 4 Set $\tilde{\alpha}^{\ell}_{y}=\tilde{\alpha}^{\ell}_{z}=(1,\ldots,1)\in\mathbb{R}^{n}$
6       $(y^{\ell+1},\tilde{\alpha}^{\ell+1}_{y})=$ DF-Search($P_{\tau_{k}}(y,x)$, $x^{\ell}$, $y^{\ell}$, $\tilde{\alpha}^{\ell}_{y}$)
7       $(z^{\ell+1},\tilde{\alpha}^{\ell+1}_{z})=$ DF-Search($P_{\tau_{k}}(z,x)$, $x^{\ell}$, $z^{\ell}$, $\tilde{\alpha}^{\ell}_{z}$)
8       Set $x^{\ell+1}=\arg\min_{x\in X}P_{\tau_{k}}(x,y^{\ell+1},z^{\ell+1})$
@@ 9      Set $\ell=\ell+1$
Clearly, the method revolves around the DF-Search subprocedure, described in Algorithm 3. In the latter, coordinate directions are explored according to classical derivative-free line search principles [20]: each direction (and its opposite) are polled for a given tentative stepsize, checking a sufficient decrease condition: if the step can be accepted, a further extrapolation is carried out as long as sufficient decrease is attained and the tentative stepsize at the subsequent iteration (for that direction) will be equal to the current actual stepsize; on the other hand, if sufficient decrease is not attained, the direction is discarded altogether for the iteration and the tentative stepsize is decreased for the next one.
1 Input: $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, $x^{\ell},w^{1}\in\mathbb{R}^{n},$ $\tilde{\alpha}^{\ell}\in\mathbb{R}^{n}$, $\theta\in(0,1)$, $\gamma\in(0,1)$
@@ 3.1 Convergence analysis
$\|\nabla_{y}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\|\leq(c_{1}+c_{2}\tau_{k})\max_{i=1,\ldots,n}\tilde{\alpha}_{yi}^{\ell+1}.$
@@
By applying the Mean Value Theorem in the above inequality in the case of direction $+d_{i}$, we have
@@
By an analogous reasoning for the case of direction $-d_{i}$ we obtain
@@
Case (b). Again, from the instructions of Algorithm 3 and assuming without loss of generality that sufficient decrease is obtained along $+d_{i}$, we have
@@
and hence it follows that the sequence $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ converges. Furthermore, by the instructions of Algorithm 3, the update of each individual variable from vectors $y$ and $z$ guarantees a sufficient decrease of $-\gamma{(\alpha_{i}^{\ell})}^{2}$ (this is trivially satisfied even when $\alpha_{i}^{\ell}=0$); we can therefore write
@@
By the instructions of Algorithm 2, $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ is a decreasing sequence, hence $\{(x^{\ell},y^{\ell},z^{\ell})\}\subseteq\mathcal{L}_{0}^{\tau_{k}}=\{(x,y,z)\mid P_{\tau_{k}}(x,y,z)\leq P_{\tau_{k}}(\hat{x}^{k-1},\hat{y}^{k-1},\hat{z}^{k-1})\}$. By Corollary 2.2, the level set $\mathcal{L}_{0}^{\tau_{k}}$ is compact, therefore the sequence $\{(x^{\ell},y^{\ell},z^{\ell})\}$ admits limit points. Let us now assume, by contradiction, that there exists an infinite subset ${T}$ such that
@@
and $\alpha_{\ell}$ is determined by the Armijo-type line search. From the properties of the Armijo-type line search ([12, Prop. 20.4]) we get
@@
By the properties of the projected gradient direction [12, Eq. (20.15)], we also have
@@ 4 The case of partially separable functions
$\min_{x\in X,y_{1},\ldots,y_{m}\in\mathbb{R}^{n}}$ $\sum_{i=1}^{m}f_{j}(y_{j})$
@@ 5 Computational experiments
We report in this section an experimental evaluation of the proposed penalty decomposition derivative free (PDDF) method. We consider a set of problems of the form (23), where the functions $f_{j}$ are partially separable, by extending some test problems from CUTEst [11] as done in [21]. Each problem is obtained taking the summation of several copies of a base function, where a subset of variables is shared between functions, following different patterns. In ARWHEAD a variable is common to all the functions in the summation. In BDEXP, ENGVAL, MOREBV and TRIDIA variables are shared between consecutive functions, whereas in DIXMAANI, DIXMAANA, NZF1 and WOODS variables are shared following a more intricate rule. In BDQRTIC we find both variables shared between consecutive functions and a variable that is shared among them all. In BEALES, POWSING and ROSENBR instead, no variable is shared between the functions, thus each term of the summation can be optimized independently. We consider different dimensionalities $n$ for each problem, ranging from 10 to 6500. In Table 1 we report in detail the dimensionalities tested for each problem and the number of functions that compose the objective function. All problems considered in this experimental analysis are unconstrained.
Table 1: Details of the problems employed in the experimental analysis.
Problem $\mathbf{n}$ $\mathbf{m}$
$(n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$
ARWHEAD 10, 50, 100, 500, 1000, 5000 $n-1$
BDEXP 10, 50, 100, 500, 1000, 5000 $n-2$
BDQRTIC 10, 50, 100, 500, 1000, 5000 $n-4$
BEALES 10, 50, 100, 500, 1000, 5000 $n/2$
BROYDN3D 10, 50, 100, 500, 1000, 5000 $n$
DIXMAANA 15, 51, 102, 501, 1002, 5001 $n$
DIXMAANI 15, 51, 102, 501, 1002, 5001 $n$
ENGVAL 10, 50, 100, 500, 1000, 5000 $n-1$
MOREBV 12, 52, 102, 502, 1002, 5002 $n$
NZF1 13, 39, 130, 650, 1300, 6500 $16n/13-$ 1
POWSING 20, 52, 100, 500, 1000, 5000 $n/4$
ROSENBR 10, 50, 100, 500, 1000, 5000 $n/2$
TRIDIA 10, 50, 100, 500, 1000, 5000 $n$
WOODS 20, 40, 200, 400, 2000, 4000 $23n/$ 2
@@
The performance of the proposed PDDF method is evaluated in comparison to the classical line-search based derivative-free coordinate descent method (LS) proposed in [20], in its unconstrained variant: objective function sampling is carried out along coordinate directions, making extrapolation steps when a sufficient decrease of the objective function is obtained for the tentative stepsize. We refer the reader to [20] for further details.

### v3

@@ Abstract
In this paper, we consider the problem of minimizing a smooth function, given as finite sum of black-box functions, over a convex set. In order to advantageously exploit the structure of the problem, for instance when the terms of the objective functions are partially separable, noisy, costly or with first-order information partially accessible, we propose a framework where the penalty decomposition approach is combined with a derivative-free line-search-based method. Under standard assumptions, we state theoretical results showing that the proposed algorithm is well-defined and globally convergent to stationary points. The results of preliminary numerical experiments, performed on test problems with number of variables up to thousands, show the validity of the proposed method w.r.t. state of the art methods.
@@
Penalty decomposition, Finite-sum, Coordinate partially separable, Derivative-free line-search
@@ 1 Introduction
$\min_{x}$ $f(x)={\sum_{j=1}^{m}}f_{j}(x)$ (1)
s.t. $x\in X$
where $X\subset\mathbb{R}^{n}$ is a nonempty, closed, convex set and $f_{j}:\mathbb{R}^{n}\to\mathbb{R}$, for $j=1,\ldots,m$ are smooth functions for which, however, we do not have access to first-order information. We also assume that $f$ is bounded below by some value $f^{*}$ and that the euclidean projection operator onto $X$ is effectively available. The latter assumption would in principle hold for any convex set; yet, while projection can be exactly and cheaply obtained for important constraints sets such as bounds or spheres, in other cases it is impractical to handle.
The literature on derivative-free methods is wide and increasing. For insights on general derivative-free approaches we refer the reader to the books [8, 2], the survey [21] and to some seminal papers both for smooth optimization (e.g., [10, 31, 22, 30]) and nonsmooth optimization (e.g., [1, 13, 23]). A very recent frontier (investigated in [6] and references therein) is that of derivative-free stochastic optimization, where the black-box provides evaluations of a stochastic objective function.
According to the classification of derivative-free methods stated in [21], we set our work within the class of direct search methods, i.e., methods that use only function values without approximating the gradient. This specific class of method is also the subject of the thorough survey [12]. More specifically, we deal with deterministic objective functions, and the proposed derivative-free approach fits that of the so-called directional direct-search methods, where the current iterate $x_{k}$ is updated by a term of the form $\alpha_{k}d_{k}$, where $\alpha_{k}$ is the step size and $d_{k}$ is a vector taken from a finite set of search directions $D_{k}$ that surely includes a feasible descent direction, so that (sufficiently) small step sizes along such direction lead to an improved feasible solution. Most directional direct-search methods share the choice of the set of search directions $D_{k}$ and differ in the strategies to sample the objective function along them. We adopt the line-search strategy, i.e., a new point is produced by performing a derivative free line-search that guarantees a sufficient decrease of the objective function.
@@
Briefly, the penalty decomposition approach consists in the introduction of an auxiliary vector of variables - copies in principle of the original ones - and the addition of a (sequential) penalty term to the objective function, penalizing the distance between each pair of corresponding variables. The introduction of the auxiliary variables allows the decomposition of the problem into smaller independent subproblems; this may be advantageously exploited to take into account the peculiarities of problem (1) in the cases (i)-(iv) discussed above.
The penalty decomposition approach is related to that of Alternating Direction Method of Multipliers (ADMM) (see, e.g., [28]), which could be used to solve the reformulated problem in place of the sequential penalty method. ADMM methods come with a strong convergence theory and perform well in practice even in some nonconvex cases, but require to compute the global minimizers of augmented Lagrangian functions. Therefore, the development and application of ADMM-based methods may be prohibitive in a general nonconvex and derivative-free setting as that of the present work where, to take into account these issues, we adopt the sequential penalty approach.
@@
In this work we define a general algorithm model where the penalty decomposition strategy is coupled with a derivative-free line-search-based method for the management of subproblems. We state theoretical results - not immediately descending from known results in the literature - for proving global convergence properties of the proposed algorithm and we present its characterization in the specific case of partially separable functions. Finally, we report the results of preliminary computational experiments.
@@ 2 A Penalty decomposition approach
A sequential penalty method [4, Ch. 4.2] can then be employed to address problem (3) and thus (2), which iteratively solves (inexactly) subproblems of the form
@@
$\text{s.t. }x\in X,$
for a sequence of values $\{\tau_{k}\}$ such that $\tau_{k}\to\infty$. We shall highlight here that, at this stage, the inner solver that tackles the subproblems is an unspecified oracle, that could possibly employ first-order (ore even further) information. The black-box nature of the proposed overall scheme will be in the following sections.
The subproblem is considered (approximately) solved if a solution $(x^{k},y^{k},z^{k})$ is found such that
@@
for some positive tolerance $\xi_{k}>0$ and constants $c_{y},c_{z}>0$. The resulting algorithm is formally described by Algorithm 1. Note that, formally, the solution $(x^{k-1},y^{k-1},z^{k-1})$ is not used during iteration $k$; however, it is well-known that, with sequential penalty type methods, using the solution of the previous subproblem as a starting point for the next one is often good for efficiency and stability reasons [4, Ch. 4.2]. We shall also observe that, in addition to the approximate optimality conditions, a further boundedness property is required for solution $(x^{k},y^{k},z^{k})$.
Remark 1.
Note that, for any $\tau>0$ and any feasible $(x,y,z)$, we have
$P_{\tau}(x,y,z)=f(x).$
Therefore, if we start with a solution $(x^{0},y^{0},z^{0})$ such that $x^{0}\in X$ and $y^{0}=z^{0}=x^{0}$, condition $P_{\tau_{k}}(x^{k},y^{k},z^{k})\leq P_{\tau_{0}}(x^{0},y^{0},z^{0})$ can certainly be satisfied if at iteration $k$ we generate $(x^{k},y^{k},z^{k})$ by means of a monotone descent method starting at
We now turn to the convergence analysis of the sequential penalty framework.
1Input: $x^{0},y^{0},z^{0}\in\mathbb{R}^{n},$ $\{\xi_{k}\}\subseteq\mathbb{R}^{+},\{\tau_{k}\}\subseteq\mathbb{R}^{+}$, $c_{y},c_{z}>0$
@@ 2.1 Convergence analysis
We start the analysis stating the following assumption (formally given for the case of $f$ defined as the sum of $m$ terms).
@@
$\mathcal{L}=\{(x,y,z)\in X\times\mathbb{R}^{n}\times\mathbb{R}^{n}\mid P_{\tau}(x,y,z)\leq P_{\tau}(\bar{x},\bar{y},\bar{z})\}$
@@
Let $\{(x^{k},y^{k},z^{k})\}$ be a sequence satisfying $P_{\tau_{k}}(x^{k},y^{k},z^{k})\leq P_{\tau_{0}}(x^{0},y^{0},z^{0})$ for all $k$. Then $\{(x^{k},y^{k},z^{k})\}$ is a bounded sequence.
@@
is compact. The instructions of the algorithm imply that
@@
The result in Proposition 2.4 cannot be directly inferred, to the best of our knowledge, by known results in the literature. Indeed, most known results concerning penalty and augmented Lagrangian methods rely on different stationarity conditions for subproblems with lower-level constraints to be handled explicitly out of the penalty function (see, e.g., [5, 7]). On the other hand, the case of projection-based stationarity has been previously considered in [14, 15], but a bond between the sequences $\{\tau_{k}\}$ and $\{\xi_{k}\}$ is required ($\tau_{k}\xi_{k}\to 0$). For this particular setup we thus provide a novel, improved result of convergence that does not need this particular interaction between the two sequences of parameters.
@@ 3 A derivative-free alternate minimization algorithm for penalty subproblems
    fixed $y$ and $z$, we have a problem in $x$ whose unique globally optimal solution can be determined, being it strictly convex quadratic with a separable objective function subject to convex constraints; the solution is in fact readily available by projecting the centroid of $y^{\ell+1}$ and $z^{\ell+1}$ (and the other blocks of variables whenever $m>2$) onto the convex set $X$.
@@ 4 Set $\tilde{\alpha}^{\ell}_{y}=\tilde{\alpha}^{\ell}_{z}=(1,\ldots,1)\in\mathbb{R}^{n}$
[…]
8   Set $x^{\ell+1}\in\arg\min_{x\in X}P_{\tau_{k}}(x,y^{\ell+1},z^{\ell+1})$
@@ 9  Set $\ell=\ell+1$
Clearly, the method revolves around the DF-Search subprocedure, described in Algorithm 3. In the latter, coordinate directions are explored according to classical derivative-free line-search principles [26]: each direction (and its opposite) are polled for a given tentative stepsize, checking a sufficient decrease condition: if the step can be accepted, a further extrapolation is carried out as long as sufficient decrease is attained and the tentative stepsize at the subsequent iteration (for that direction) will be equal to the current actual stepsize; on the other hand, if sufficient decrease is not attained, the direction is discarded altogether for the iteration and the tentative stepsize is decreased for the next one.
1Input: $g:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$, $x^{\ell},w^{1}\in\mathbb{R}^{n},$ $\tilde{\alpha}^{\ell}\in\mathbb{R}^{n}$, $\theta\in(0,1)$, $\gamma\in(0,1)$
@@ 3.1 Convergence analysis
$\|\nabla_{y}P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\|\leq(c_{1}+c_{2}\tau_{k})\max_{i=1,\ldots,n}\tilde{\alpha}_{yi}^{\ell+1},$
where $\hat{\alpha}_{yi}^{\ell+1}$ denotes $(\tilde{\alpha}_{y}^{\ell+1})_{i}$.
@@
By applying the Mean Value Theorem in the above inequality in the case of direction $d_{i}$, we have
@@
By an analogous reasoning for the case of direction $d_{i}$ we obtain
@@
Case (b). Again, from the instructions of Algorithm 3 and assuming without loss of generality that sufficient decrease is obtained along $d_{i}$, we have
@@
and hence it follows that the sequence $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ converges. Furthermore, by the instructions of Algorithm 3, the update of each individual variable from vectors $y$ and $z$ guarantees a sufficient decrease of $\gamma{(\alpha_{i}^{\ell})}^{2}$ (this is trivially satisfied even when $\alpha_{i}^{\ell}=0$); we can therefore write
@@
By the instructions of Algorithm 2, $\{P_{\tau_{k}}(x^{\ell},y^{\ell},z^{\ell})\}$ is a decreasing sequence, hence $\{(x^{\ell},y^{\ell},z^{\ell})\}\subseteq\mathcal{L}_{0}^{\tau_{k}}=\{(x,y,z)\mid P_{\tau_{k}}(x,y,z)\leq P_{\tau_{k}}(\hat{x}^{k-1},\hat{y}^{k-1},\hat{z}^{k-1})\}$. By Corollary 2.2, the level set $\mathcal{L}_{0}^{\tau_{k}}$ is compact, therefore the sequence $\{(x^{\ell},y^{\ell},z^{\ell})\}$ admits limit points. Let us now assume, by contradiction, that there exists an infinite subset T{T} such that
@@
and $\alpha_{\ell}$ is determined by the Armijo-type line-search. From the properties of the Armijo-type line-search ([18, Prop. 20.4]) we get
@@
By the properties of the projected gradient direction [18, Eq. (20.15)], we also have
@@ 4 The case of partially separable functions
$\min_{\begin{subarray}{c}x\in X,\\ y_{j}\in\mathbb{R}^{|S_{j}|}\end{subarray}}$ $\sum_{j=1}^{m}f_{j}(y_{j})$
@@ 5 Computational experiments
We report in this section an experimental evaluation of the proposed penalty decomposition derivative free (PDDF) method. We consider a set of problems of the form (22), where the functions $f_{j}$ are partially separable, by extending some test problems from CUTEst [16]. Similarly to what is done in [29], we obtain the test problems by taking the summation of several copies of base functions, where a subset of variables is shared between functions, following different patterns. In ARWHEAD a variable is common to all the functions in the summation. In BROWNAL6, BROYDN3D, ENGVAL, FREUROTH, MOREBV and TRIDIA variables are shared between consecutive functions, whereas in DIXMAANA, DIXMAANI, NZF1 and WOODS variables are shared following a more intricate rule. In BDQRTIC we find both variables shared between consecutive functions and a variable that is shared among them all. In BEALES, POWSING and ROSENBR instead, no variable is shared between the functions, thus each term of the summation can be optimized independently. The other problems listed in Table 1 are constructed by taking $m$ copies of each base CUTEst problem and linking them sequentially, so that the last variable of one copy is shared with the first variable of the next one. Additional problems have been generated considering the summation of different CUTEst functions and sharing the first variables between those. The details of the problems constructed in this latter way are available in Table 2. The base CUTEst problems that we used in this work comprise academic, modeling and real-world problems, both unconstrained and with box constraints. We consider different dimensionalities $n$ for the problems, ranging from 4 to 1000. In total we have 231 unconstrained problems and 44 constrained ones.
Table 1: Details of the problems employed in the experimental analysis. The problems are obtained taking the summation of several copies of the base CUTEst problem reported and sharing a subset of variables between functions.The problem name also reports whether the problem is unconstrained [UC] or box-constrained [B].
Problem $\mathbf{n}$ $\mathbf{m}$ Problem $\mathbf{n}$ $\mathbf{m}$
AKIVA [UC] 5, 9, 17, 33 4, 8, 16, 32 HATFLDGLS [UC] 97, 193, 385, 769 4, 8, 16, 32
ARWHEAD [UC] 10, 50, 100, 500, 1000 9, 49, 99, 499, 999 HILBERTA [UC] 5, 9, 17, 33 4, 8, 16, 32
BA-L1SPLS [UC] 225, 449, 897 4, 8, 16 HILBERTB [UC] 37, 73, 145, 289 4, 8, 16, 32
BARD [UC] 9, 17, 33, 65 4, 8, 16, 32 HIMMELBCLS [UC] 5, 9, 17, 33 4, 8, 16, 32
BDQRTIC [UC] 10, 50, 100, 500, 1000 6, 46, 96, 496, 996 HIMMELBG [UC] 5, 9, 17, 33 4, 8, 16, 32
BEALE [UC] 5, 9, 17, 33 4, 8, 16, 32 HS110 [B] 37, 73, 145, 289 4, 8, 16, 32
BEALES [UC] 10, 50, 100, 500, 1000 5, 25, 50, 250, 500 LRIJCNN1 [UC] 85, 169, 337, 673 4, 8, 16, 32
BIGGS6 [UC] 21, 41, 81, 161 4, 8, 16, 32 LUKSAN11LS [UC] 397, 793 4, 8
BOXBODLS [UC] 5, 9, 17, 33 4, 8, 16, 32 LUKSAN12LS [UC] 389, 777 4, 8
BRKMCC [UC] 5, 9, 17, 33 4, 8, 16, 32 LUKSAN13LS [UC] 389, 777 4, 8
BROWNAL6 [UC] 10, 50, 102, 502 2, 12, 25, 125 LUKSAN14LS [UC] 389, 777 4, 8
BROYDN3D [UC] 10, 50, 100, 500, 1000 10, 50, 100, 500, 1000 LUKSAN21LS [UC] 397, 793 4, 8
CHNROSNB [UC] 197, 393, 785 4, 8, 16 MOREBV [UC] 12, 52, 102, 502 12, 52, 102, 502
CHNRSNBM [UC] 197, 393, 785 4, 8, 16 NZF1 [UC] 13, 39, 130, 650 5, 17, 59, 299
CLUSTERLS [UC] 5, 9, 17, 33 4, 8, 16, 32 POWERSUM [UC] 13, 25, 49, 97 4, 8, 16, 32
COOLHANSLS [UC] 33, 65, 129, 257 4, 8, 16, 32 POWERSUMB [B] 13, 25, 49, 97 4, 8, 16, 32
DENSCHNA [UC] 5, 9, 17, 33 4, 8, 16, 32 POWSING [UC] 20, 52, 100, 500 5, 13, 25, 125
DENSCHNB [UC] 5, 9, 17, 33 4, 8, 16, 32 QING [UC] 397, 793 4, 8
DENSCHND [UC] 9, 17, 33, 65 4, 8, 16, 32 QINGB [B] 17, 33, 65, 129 4, 8, 16, 32
DEVGLA1 [UC] 13, 25, 49, 97 4, 8, 16, 32 ROSENBR [UC] 10, 50, 100, 500, 1000 5, 25, 50, 250, 500
DEVGLA1B [B] 13, 25, 49, 97 4, 8, 16, 32 SNAIL [UC] 5, 9, 17, 33 4, 8, 16, 32
DGOSPEC [B] 9, 17, 33, 65 4, 8, 16, 32 SSI [UC] 9, 17, 33, 65 4, 8, 16, 32
DIAGIQB [B] 37, 73, 145, 289 4, 8, 16, 32 TESTQUAD [UC] 37, 73, 145, 289 4, 8, 16, 32
DIXMAANA [UC] 15, 51, 102, 501 15, 51, 102, 501 TOINTGOR [UC] 197, 393, 785 4, 8, 16
DIXMAANI [UC] 15, 51, 102, 501 15, 51, 102, 501 TOINTPSP [UC] 197, 393, 785 4, 8, 16
ENGVAL [UC] 10, 50, 100, 500, 1000 9, 49, 99, 499, 999 TOINTQOR [UC] 197, 393, 785 4, 8, 16
ERRINROS [UC] 197, 393, 785 4, 8, 16 TRIDIA [UC] 10, 50, 100, 500, 1000 10, 50, 100, 500, 1000
ERRINRSM [UC] 197, 393, 785 4, 8, 16 TRIGON1 [UC] 37, 73, 145, 289 4, 8, 16, 32
EXPFIT [UC] 5, 9, 17, 33 4, 8, 16, 32 TRIGON1B [B] 37, 73, 145, 289 4, 8, 16, 32
FREUROTH [UC] 10, 50, 100, 500, 1000 9, 49, 99, 499, 999 VANDANMSLS [UC] 85, 169, 337, 673 4, 8, 16, 32
HATFLDA [B] 13, 25, 49, 97 4, 8, 16, 32 WATSON [UC] 45, 89, 177, 353 4, 8, 16, 32
HATFLDB [B] 13, 25, 49, 97 4, 8, 16, 32 WOODS [UC] 20, 40, 200, 400 30, 60, 300, 600
HATFLDFLS [UC] 9, 17, 33, 65 4, 8, 16, 32
Table 2: Details of the problems employed in the experimental analysis. The problems are obtained taking the summation of different CUTEst problems and sharing the first $s$ variables between functions. The problem name also reports whether the problem is unconstrained [UC] or box-constrained [B].
Problem $\mathbf{n}$ $\mathbf{m}$ $\mathbf{s}$
3PK, CHEBYQAD, DECONVB, SANTALS [B] 178 4 8
AKIVA, BEALE, BOXBODLS, BRKMCC, BROWNBS [UC] 6 5 1
ALLINITU, BROWNDEN, BIGGS6, FBRAIN3LS, STREG [UC] 16 5 2
BA-L1LS, BA-L1SPLS [UC] 106 2 8
BARD, BOX3, CLIFF, CLUSTERLS, CUBE, DJTL, ENGVAL2, EXPFIT, GAUSSIAN, GROWTHLS, HAIRY [UC] 17 11 1
BQPGABIM, BQPGASIM [B] 88 2 8
CHNROSNB, CHNRSNBM [UC] 92 2 8
CHWIRUT1LS, CHWIRUT2LS [UC] 4 2 2
DENSCHNA, DENSCHNB, DENSCHNC, DENSCHND, DENSCHNE, DENSCHNF [UC] 9 6 1
DEVGLA1B, DEVGLA2B [B] 7 2 2
DIAMON2DLS, DIAMON3DLS [UC] 157 2 8
FBRAIN2LS, FBRAINLS [B] 4 2 2
GAUSS1LS, GAUSS2LS, GAUSS3LS [UC] 20 3 2
HATFLDA, HATFLDB, HATFLDC [B] 25 3 4
HEART6LS, HEART8LS [UC] 12 2 2
HS1, HS110, HS2, HS25, HS3, HS38, HS3MOD, HS4, HS5 [B] 21 9 1
HYDC20LS, HYDCAR6LS [UC] 112 2 16
LANCZOS1LS, LANCZOS2LS, LANCZOS3LS [UC] 14 3 2
LUKSAN11LS, LUKSAN12LS, LUKSAN13LS, LUKSAN14LS, LUKSAN15LS, LUKSAN16LS, LUKSAN17LS, LUKSAN21LS, LUKSAN22LS [UC] 766 9 16
MANCINO, QING, SENSORS [UC] 236 3 32
MGH09LS, MGH10LS, MGH10SLS, MGH17LS, MGH17SLS, NELSONLS, OSBORNEA, OSBORNEB [UC] 25 8 2
MISRA1ALS, MISRA1BLS, MISRA1CLS, MISRA1DLS [UC] 5 4 1
PFIT1LS, PFIT2LS, PFIT3LS, PFIT4LS [B] 6 4 2
TOINTGOR, TOINTPSP, TOINTQOR [UC] 134 3 8
TRIGON1, TRIGON2 [UC] 18 2 2
TRIGON1B, TRIGON2B [B] 16 2 4
VESUVIALS, VESUVIOLS, VESUVIOULS [UC] 20 3 2
We also consider one problem outside of CUTEst collection, that is the Lockwood problem [17]. In its original formulation, we aim at minimizing a linear objective with two nonlinear constraints and box constraints, and is formulated as
$\min_{x}\left\{f(x)=\sum_{j=1}^{6}x_{j}:c_{1}(x)\leq 0,c_{2}(x)\leq 0,x\in[0,2\cdot 10^{4}]^{6}\right\}.$ (23)
Since the two nonlinear constraints are black boxes, we are not able to use PDDF to solve the Lockwood problem directly. We therefore consider two alternative versions of this problem, that consider the two nonlinear constraints as part of the objective, i.e., in the form of penalties. We will elaborate on this later.
@@
The performance of the proposed PDDF method is evaluated in comparison to the classical line-search based derivative-free coordinate descent method (LS) proposed in [26]. In LS the objective function sampling is carried out along coordinate directions, making extrapolation steps when a sufficient decrease of the objective function is obtained for the tentative stepsize, ensuring at every step that the solution remains feasible by using a maximal step size. We refer the reader to [26] for further details. In addition, we compare our algorithm with the NOMAD implementation [3] of the Mesh Adaptive Direct Search algorithm, a widely used derivative-free optimization method that explores the search space using a set of mesh points whose resolution is adaptively refined to efficiently locate solutions.